[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO Biweekly Status Report

  Compton Observatory Science Report #170, Monday November 14, 1994
      Chris Shrader, Compton Observatory Science Support Center

         Questions or comments can be sent to the CGRO SSC.
          Phone: 301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                          Spacecraft Status

The spacecraft and all subsystems continue to function normally. A
maneuver was uplinked on November 9, 1994 to point the Z-axis towards
a target of opportunity, OJ 287, a BL Lac object which is undergoing
significant photometric activity. These observations are being made
under a Guest Investigator Program, and are part of a multiwavelength

As noted in the EGRET instrument report below, the EGRET has been
operated in a mode which reduces the field of view, and thus the
number of spark triggers. It is anticipated that this operating mode
will be used extensively in Cycle 5 and beyond to prolong the
lifetime of the spark-chamber gas.

A new verision of the Cycle-4 timeline, which reflects the changes
resulting from the OJ 287 ToO has been sent to all Guest
Investigators, and is also available on GRONEWS
(grossc.gsfc.nasa.gov, username = gronews) or Xmosiac

                         Instrument Reports

EGRET operations were normal this biweekly period. Delivery of data
to the GRO SSC remains on schedule. Interaction with guest
investigators continues at a good level.

For period 403, EGRET was in the narrow telescopes only mode. When
the target of opportunity mode began on Wednesday November 9, 1994,
at the request of the Project Scientist and with our concurrence, it
was left in the narrow telescopes only mode. It will be placed in the
normal mode for viewing period 404.

Several papers will be submitted to journals soon summarizing EGRET
results for the entire sky. There will also be several talks
presented at the Munich Relativistic Astrophysics meeting.


OSSE operations are normal. A recent modification to an address in
flight software resulted in our missing both of last week's slewable
burst triggers from BATSE. The address has been updated, and we plan
to get everything right for the next trigger.

In viewing period 403.0 (1-9 Nov), the Z-axis target was Her X-1, and
the X-axis targets were IC4329A and NGC 4945 (Guest Investigator R.
Mushotzky). In Target of Opportunity viewing period 403.5 (9-15 Nov),
the Z-axis target is the blazar OJ 287, and the X-axis targets are PG
1416-249, GRS 1716-249, and GX1+4. BATSE reports that both of the
latter objects are currently in high states.

Data from viewing periods 229, 229.3, 229.5, 230, 230.5, and 231 were
delivered to the Compton GRO Science Support Center archive this
week. The targets during these viewing periods are listed below.

     vp        targets
     229       galactic plane near (l,b)=(5,0), and Coma Cluster
     229.3     PSR 1509-58
     229.5     galactic plane near (l,b)=(5,0), and Coma Cluster
     230       GRO J1008-57, Sun, and QSO 0716+714.
     230.5     GRO J1008-57, Sun, and QSO 0716+714
     231       NGC 6814 and NGC 4507


The COMPTEL instrument is performing well and continues routine

The COMPTEL Burst Rapid Response Team achieved a new breakthrough
recently in the rapid processing of burst data. In the first test of
a new automated data-processing system for bursts, an image of the
COMPTEL field of view was obtained 8 minutes after the receipt of
data from the BACODINE system; the transfer of BACODINE data occurred
within minutes of a BATSE burst trigger on Sunday, 23 October. While
no emission from this particular burst was detected in the COMPTEL
energy range, this exercise demonstrates the practical feasibility of
reducing the COMPTEL processing time to approximately 10 to 15
minutes for burst events occurring within the field of view of the
instrument. The team will continue to further refine the data
processing procedures for bursts. Ground-based observers interested
in receiving COMPTEL burst alerts within approximately 10-15 minutes
of burst occurrence should contact Dr. Bernie McNamara (NMSU) at

A general team meeting of the collaboration is being held this week
at the Max-Planck-Institute in Garching, Germany.


A new outburst of the x-ray transient GRO J1655-40 began on November
1st. By November 11th and 12th the 20-100 kev flux had increased to
nearly 1 Crab. The x-ray transients GX 339-4 and GRO J1719-24 = GRS
1716-249 continue to be in outburst.

The following sources were detected by the BATSE pulsed source
monitor in the past two weeks: Her X-1, Cen X-3, 4U 1626-67, 2S
1417-624, GX 1+4, Vela X-1, and GX 301-2. The outburst of 2S 1417-624
has now lasted 81 days. The barycentric pulse frequency is again stop
increasing, possibly indicating a second periastron passage. The
x-ray binary GX 1+4, which has been spinning down since 1987, has
been spinning up for the last week.

As of November 12th, BATSE has detected 1158 cosmic gamma-ray bursts
out of a total of 3177 on-board triggers in 1299 days of operation.
There have been 740 triggers due to solar flares with emission above
60 keV. Since the BATSE burst trigger criteria were changed on
September 19th, there have been 7 triggers due to atmospheric
gamma-ray flashes.